A person scores 81 on a test of verbal ability and 6.4 on a test of quantitative
ability. For the verbal ability test, the mean for people in general is 50 and the
standard deviation is 20. For the quantitative ability test, the mean for people in
general is 0 and the standard deviation is 5. Which is this person’s stronger ability:
verbal or quantitative? Explain your answer to a person who has never had a
course in statistics.

Answer by edjones(8007)

z=(81-50)/20=1.55
p-value of 1.55 is 0.939 so this person did better than 93.9% of those taking the test
.
z=(6.4-0)/5=1.25
p-value of 1.25 is 0.894 so this person did better than 89.4% of those taking the test
.
Ed